Silaoreo123:

The distance from earth to the North Star is 4.08 x10^12km . What is this number written in its long form?

jaycox:

You have to use a calculator and put in the problem

jhonyy9:

look please one example 5*10^2 = 5*100 = 500 so using this example can you solve correct your posted problem ?